The velocity of a train is reduced uniformly from 40km sec to 30kmsec while travelling a distance of 100km.


a=calculate the retardation


b=how much farther will it travel before coming to rest, assuming the same retardation.

1
Expert's answer
2021-11-11T17:09:19-0500

a) "s=\frac{v_f^2-v_i^2}{2a}\to a=\frac{v_f^2-v_i^2}{2s}=\frac{30^2-40^2}{2\cdot100}=-3.5\ (km/s^2)"


b) "s=\frac{v_f^2-v_i^2}{2a}=\frac{0^2-30^2}{2\cdot (-3.5)}=128.57\ (km)"